OpenStudy (anonymous):

Let log[b] A=1, log[b] C=3, log[b] D=4 What is the value of log[b] (A^5 C^2/D^6) Im confused can someone explain how to do this? A. 26 B. 0 C. There isn't enough information to answer the question. D. -13 How is -13 the answer ????

OpenStudy (anonymous):

\[5\log _{b}A + 2\log _{b}C- 6\log _{b}D\]

OpenStudy (amistre64):

seems fair

OpenStudy (anonymous):

but how would it become -13?

OpenStudy (amistre64):

\[5\log _{b}A + 2\log _{b}C- 6\log _{b}D\] but they tell you the log parts, sub them in \[5(1)+ 2(3)- 6(4)\]

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhh! 5+6-24 = -13 !!!! ohhhhh thank youuuu @amistre64

OpenStudy (amistre64):

:) yep

OpenStudy (anonymous):

i didnt see that lol i was confuse cuz 5+2-6 was definitely not -13 thank you @amistre64